OpenStudy (anonymous):

alternating series of (-1)^n (ln(n))/n^2 can i compare the series to a 1/n^2 convergent series and then say that by limit comparison test the series absolutely converges?

OpenStudy (nottim):

What subject is this?

OpenStudy (anonymous):

math...

OpenStudy (nottim):

wow. then I must be super dumb.

OpenStudy (anonymous):

? amm is calculus 2 ... alternating series.. ... .. amm i think im still doing math lol

OpenStudy (zarkon):

\[\ln(n)\le \sqrt{n}\]

OpenStudy (zarkon):

\[\frac{\ln(n)}{n^2}\le\frac{\sqrt{n}}{n^2}=\frac{1}{n^{3/2}}\]

OpenStudy (experimentx):

Yup, Zarkon is right, 1/n^(3/2) converges so left hand side must also converge. we can show that (-1)^n * (ln(n))/n^2 converges if (ln(n))/n^2 converges, since absolute value is greater than the other alternating value

OpenStudy (anonymous):

ok but why compareto to ((n)^1/2))/(n^2)???

OpenStudy (zarkon):

why not...it works

OpenStudy (experimentx):

because ((n)^1/2))/(n^2) is a convergent series and your series is less than it

OpenStudy (anonymous):

got it thanks guys...